$SPAD/src/input symbol.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool symbol.output
)set message test on
)set message auto off
)clear all
--S 1 of 24
X: Symbol := 'x
--R
--R
--R   (1)  x
--R                                                                 Type: Symbol
--E 1
--S 2 of 24
XX: Symbol := x
--R
--R
--R   (2)  x
--R                                                                 Type: Symbol
--E 2
--S 3 of 24
A := 'a
--R
--R
--R   (3)  a
--R                                                            Type: Variable(a)
--E 3
--S 4 of 24
B := b
--R
--R
--R   (4)  b
--R                                                            Type: Variable(b)
--E 4
--S 5 of 24
x**2 + 1
--R

--R
--R         2
--R   (5)  x  + 1
--R                                                    Type: Polynomial(Integer)
--E 5
--S 6 of 24
"Hello"::Symbol
--R
--R
--R   (6)  Hello
--R                                                                 Type: Symbol
--E 6
--S 7 of 24
new()$Symbol
--R
--R
--R   (7)  %A
--R                                                                 Type: Symbol
--E 7
--S 8 of 24
new()$Symbol
--R
--R
--R   (8)  %B
--R                                                                 Type: Symbol
--E 8
--S 9 of 24
new("xyz")$Symbol
--R
--R
--R   (9)  %xyz0
--R                                                                 Type: Symbol
--E 9
--S 10 of 24
X[i,j]
--R
--R
--R   (10)  x
--R          i,j
--R                                                                 Type: Symbol
--E 10
--S 11 of 24
U := subscript(u, [1,2,1,2])
--R

--R
--R   (11)  u
--R          1,2,1,2
--R                                                                 Type: Symbol
--E 11
--S 12 of 24
V := superscript(v, [n])
--R
--R
--R          n
--R   (12)  v
--R                                                                 Type: Symbol
--E 12
--S 13 of 24
P := argscript(p, [t])
--R
--R
--R   (13)  p(t)
--R                                                                 Type: Symbol
--E 13
--S 14 of 24
scripted? U
--R
--R
--R   (14)  true
--R                                                                Type: Boolean
--E 14
--S 15 of 24
scripted? X
--R
--R
--R   (15)  false
--R                                                                Type: Boolean
--E 15
--S 16 of 24
string X
--R
--R
--R   (16)  "x"
--R                                                                 Type: String
--E 16
--S 17 of 24
name U
--R

--R
--R   (17)  u
--R                                                                 Type: Symbol
--E 17
--S 18 of 24
scripts U
--R
--R
--R   (18)  [sub= [1,2,1,2],sup= [],presup= [],presub= [],args= []]
--IType: Record(sub: List(OutputForm),sup: List(OutputForm),...
--E 18
--S 19 of 24
name X
--R
--R
--R   (19)  x
--R                                                                 Type: Symbol
--E 19
--S 20 of 24
scripts X
--R
--R
--R   (20)  [sub= [],sup= [],presup= [],presub= [],args= []]
--IType: Record(sub: List(OutputForm),sup: List(OutputForm),...
--E 20
--S 21 of 24
M := script(Mammoth, [[i,j],[k,l],[0,1],[2],[u,v,w]])
--R
--R
--R         0,1       k,l
--R   (21)     Mammoth   (u,v,w)
--R           2       i,j
--R                                                                 Type: Symbol
--E 21
--S 22 of 24
scripts M
--R
--R
--R   (22)  [sub= [i,j],sup= [k,l],presup= [0,1],presub= [2],args= [u,v,w]]
--IType: Record(sub: List(OutputForm),sup: List(OutputForm),...
--E 22
--S 23 of 24
N := script(Nut, [[i,j],[k,l],[0,1]])
--R

--R
--R         0,1   k,l
--R   (23)     Nut
--R               i,j
--R                                                                 Type: Symbol
--E 23
--S 24 of 24
scripts N
--R
--R
--R   (24)  [sub= [i,j],sup= [k,l],presup= [0,1],presub= [],args= []]
--IType: Record(sub: List(OutputForm),sup: List(OutputForm),...
--E 24
)spool
)lisp (bye)

———-

References^{[1] nothing}

8